what is: y= x+7 (eq. 1)
y= -4x-3 (eq. 2)

Answer by jigs_ipan(5)

equate the first and second equation.
x+7=-4x-3
x+4x=-3-7 (combine liked terms)
5x=-10 (divide both sides by 5)
x=-2
subtitute the value of x to any of the two equation to find y.
y=x+7 (eq. 1)
y=(-2)+7 (subtituting the value of x)
y=5